Problem 14) : An enhanced Voituge Regulator with a BTT Current Booster. By crocling to it a power BJT device as shown,
Answer #1

Solution:

Answer i Giren a Power BJT device. ALCE - 03A INPUT A RIPPLE IB sa 4123to. 184 DZ Dc Source. DZ Vzro (22? ß = 49 LIBE(On) -0.(a) By what factor is the De output Current Capability of the coltage regulator being increased through the use of ColoentLCE=184 -12.3 - 5.76. 12 = 12.3+UBE V2 = 12.3.to.7=136. TE = 0.34 =CB+1) IB FB FOR 50.6 mt. Ibt [2 = 18 12 18 13 Sis Flomat .Go-to tale dere Som slang talker. Gy= b = 1.39m = 1912 = 1.2m GL = 1213 - 19 GL=244m o 5m 5m f1.7m +2440412 240.8.140 upr - T(d) Determine coltage Source the abole same ckt is taken win is grounded. fit 2 et = 44 (z tamtho] Rip = L = ab tgimthe Rim -DC input power = 18 x 3in De output power = 18 [lomto.3] ud XIE 512.3 X0.3 12.3X0.3 7 Efficiency = -xloo 18 1oth to:3] :. l.E